# Discrete Approximate Circle Bundles

### v2

@@ A.3 Miscellaneous
Suppose $V=\{v_{j}\}_{j=0}^{n}\subset\mathbb{S}^{1}\subset\mathbb{R}^{2}$ such that $|v_{j}-v_{0}|<\sqrt{2}$ for all $j$. Then $V$ has a unique weighted Karcher mean with respect to the geodesic distance on $\mathbb{S}^{1}$ (and any weights).
@@
The fact that $|v_{j}-v_{0}|<\sqrt{2}$ for all $j$ implies $d_{\mathbb{S}^{1}}(v_{j},v_{0})<\frac{\pi}{2}$ for all $j$, so by the triangle inequality, the diameter of $V$ (with respect to $d_{\mathbb{S}^{1}}$) is less than $\pi$. Since $\mathbb{S}^{1}$ is geodesically convex on any subset of diameter less than $\pi$, $V$ has a unique weighted Karcher mean with respect to any weights.

### v3

@@ A.3 Miscellaneous
Here we provide several miscellaneous results which are used throughout the paper.
@@
Suppose $V=\{v_{j}\}_{j=0}^{n}\subset\mathbb{S}^{1}\subset\mathbb{R}^{2}$ such that $|v_{j}-v_{0}|<\sqrt{2}$ for all $j$. Then $V$ has a unique weighted Karcher mean with respect to the geodesic distance on $\mathbb{S}^{1}$ and any weights. In particular, given weights $w_{0},...,w_{n}$, the unique Karcher mean is given by
$\mu=\exp\left(\sum_{j=0}^{n}w_{j}\log(v_{j})\right)$
where $\log$ denotes any branch of $\log$ whose domain contains $V$.
@@
The fact that $|v_{j}-v_{0}|<\sqrt{2}$ for all $j$ implies $d_{\mathbb{S}^{1}}(v_{j},v_{0})<\frac{\pi}{2}$ for all $j$, so by the triangle inequality, the diameter of $V$ (with respect to $d_{\mathbb{S}^{1}}$) is less than $\pi$. Since $\mathbb{S}^{1}$ is geodesically convex on any subset of diameter less than $\pi$, $V$ has a unique weighted Karcher mean with respect to any weights. In particular, we can choose lifts $\{\phi_{j}\}_{j=0}^{n}\subset\mathbb{R}$ of $\{v_{j}\}_{j=0}^{n}$ which lie in an interval $I\subset\mathbb{R}$ of length less than $\frac{1}{2}$. The Frechet variance of $V$ around any $p=e^{i\phi_{p}}\in\exp(I)\subset\mathbb{S}^{1}$ is given by
$F_{V}(p)$ $=\sum_{j=0}^{n}w_{j}d_{\mathbb{S}^{1}}(p,v_{j})^{2}$
$=\sum_{j=0}^{n}w_{j}d_{\mathbb{S}^{1}}(e^{i\phi_{p}},e^{i\phi_{j}})^{2}$
$=\sum_{j=0}^{n}w_{j}d_{\mathbb{R}}(\phi_{p},\phi_{j})^{2}$
which is the weighted variance of $\{\phi_{j}\}_{j=0}^{n}$ about $\phi_{p}$. This is uniquely minimized at $\phi_{p}=\sum_{j=0}^{n}w_{j}\phi_{j}$, so $F_{V}$ has a minimizer at $p=\exp(\sum_{j=0}^{n}w_{j}\phi_{j})$. Since we have shown that weighted Karcher mean is unique under the hypotheses, $p$ is the unique weighted Karcher mean of $V$.